[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Feb 28, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Feb 28 14:48:18 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The ballyhoo N-AWIPS workstation did not come back after a reboot yesterday  
afternoon. NCO had to reload the video drivers, which solved the problem.


HSU:
Mike Brennan gave the NCO visitors a demonstration of the problems with  
reloading color tables on in N-AWIPS and they now better appreciate the  
problem. They will continue to work towards a RHEL6 configuration that has  
performance similar to the current RHEL5 configuration.
Other items:
We lost the GOES satellite ingest yesterday afternoon due to modifications  
by remote NCO staff to our satellite servers. NCO had to roll back the  
changes and the data was restored after about an hour. New rules were put  
in place so that TSB and the local NCO staff are notified of any changes to  
our satellite processing system, even those that are expected to  
be "transparent".
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170228/a78c1288/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list